Question

Ammonia is made from a rxn of nitrogen and hydrogen:

N2(g)+3H2(g)--> 2NH3

a) What mass of the ammonia is made from 2000g of N2 and 1000 g of H2?

b) What is the mass of the excess reactant?
Answer #1

N2(g)+3H2(g)--> 2NH3

moles of N2 = 2000g/ 28= 71.43

moles of H2 = 1000g/2=500

1mole of N2 react with 3 mole of H2 to form 2 mole of NH3

so, 71.43 moles of N2 react with 3*71.43 =214.29 mole of H2 to form 2*71.43 mole of NH3=142.86

a) mass of ammonia made = 142.86*17 =2428.62g

b) mass of excess reactant = (500-214.29) * 2 = 571.42g